Terrell drove to the mountains at an average of 60 miles per hour. His return trip by the same roads averaged 40 miles per hour. His total driving time was 12 hours. How far did he drive one way?
60x = 720
x = 12
60(1/3) = 240
60(2/3) = 480
: Terrell drove to the mountains at an average of 60 miles per hour. His return trip by the same roads averaged 40 miles per hour. His total driving time was 12 hours. How far did he drive one way?
60x = 720
x = 12
60(1/3) = 240
60(2/3) = 480

Answer by ankor@dixie-net.com(4538)

Terrell drove to the mountains at an average of 60 miles per hour. His return trip by the same roads averaged 40 miles per hour. His total driving time was 12 hours. How far did he drive one way?
:
let d = distance one way
:
Write a time equation: Time = dist/speed
;
To time + return time = 12 hrs
d/60 + d/40 = 12
Multiply equation by 120 to get rid of the denominators, results:
2d + 3d = 12(120)
:
5d = 1440
d = 1440/5
d = 288 mi one way:
:
:
Check solution by finding the total time
288/60 + 288/40 =
4.8 + 7.2 = 12